# Recent Event Camera Innovations: A Survey

### v1

@@ Abstract
Event cameras Event-based vision Neuromorphic vision Dynamic vision sensors Low latency High dynamic range Low power
1 Understanding Event-based Vision – An Introduction
@@ 5 Event Camera Models: An Overview
| - | - | Supports multi-camera time synchronization | - | ||||||
@@
| METAVISION Intelligence Suite - SDK support by PROPHESEE | ||||||||||||
@@ 7 Event Cameras in Action: Diverse Tasks and Impacts
| 2023 | PEDRo [17] | DAVIS346 | Persons |
@@ 8.1 Real-world Datasets
| 2024 | Event-KITTI [294] | V2E | Vehicles, pedestrians, cyclists, etc., | Yes |
| 2023 |
@@
| 2015 | N-MNIST [184] | DVS camera | 10 classes of digits (0-9) | Yes | Object Classification | An event-based version of the original MNIST dataset | ||||||||
| 2015 | N-Caltech 101 [184] |
@@
| 2015 | MNIST-DVS [229] | DVS camera | 10 classes of digits (0-9) | Yes | Object Classification | An event-based version of the original MNIST dataset |

### v2

@@ Abstract
Event cameras Event-based vision Neuromorphic vision Dynamic vision sensors Low latency High dynamic range Low power1 Understanding Event-based Vision – An Introduction
@@ 5 Event Camera Models: An Overview
- | - | Supports multi-camera time synchronization | - | ||||||
@@
METAVISION Intelligence Suite - SDK support by PROPHESEE |
@@ 7 Event Cameras in Action: Diverse Tasks and Impacts
2023 | PEDRo [17] | DAVIS346 | Persons |
@@ 8.1 Real-world Datasets
2024 | Event-KITTI [294] | V2E | Vehicles, pedestrians, cyclists, etc., | Yes |
2023 |
@@
2015 | N-MNIST [184] | DVS camera | 10 classes of digits (0-9) | Yes | Object Classification | An event-based version of the original MNIST dataset | ||||||||
2015 | N-Caltech 101 [184] |
@@
2015 | MNIST-DVS [229] | DVS camera | 10 classes of digits (0-9) | Yes | Object Classification | An event-based version of the original MNIST dataset |